2(3x-9)=6(x-3) The answer I got was no solution, is that correct.
Please advise

Answer by nerdybill(5404)

2(3x-9)=6(x-3)
Expanding both sides:
6x-18=6x-18
6x=6x
x=x
.
x = {set of all real numbers}